# A Self-Organizing Clustering System for Unsupervised Distribution Shift Detection

### v1

@@ Abstract
Modeling non-stationary data is a challenging problem in the field of continual learning, and data distribution shifts may result in negative consequences on the performance of a machine learning model. Classic learning tools are often vulnerable to perturbations of the input covariates, and are sensitive to outliers and noise, and some tools are based on rigid algebraic assumptions. Distribution shifts are frequently occurring due to changes in raw materials for production, seasonality, a different user base, or even adversarial attacks. Therefore, there is a need for more effective distribution shift detection techniques.
In this article, we introduce a self-organized continual learning framework designed for monitoring and detecting distribution changes. We explore the problem in a latent space generated by a bio-inspired self-organizing clustering and statistical aspects of the latent space. In particular, we investigate the projections made by two topology-preserving maps: the Self-Organizing Map and the Scale Invariant Map. Our method can be applied in both a supervised and an unsupervised context. We construct the assessment of changes in the data distribution as a comparison of Gaussian signals, making the proposed method fast and robust. We compare it to other unsupervised techniques, specifically Principal Component Analysis (PCA) and Kernel-PCA. Our comparison involves conducting experiments using sequences of images (based on MNIST and injected shifts with adversarial samples), chemical sensor measurements, and the environmental variable related to ozone levels. The empirical study reveals the potential of the proposed approach.
@@ 2.2 Dimensionality reduction using self-organizing clustering methods
A common approach for analyzing high-dimensional data is to apply dimensionality reduction (DR) using linear or random projections. However, both projections to the marginals and PCA present limitations for capturing the relevant structure of the original data that can cause data shift [1, 18]. On the contrary, some authors recognize the benefits of using PCA as the DR technique in a multivariate unsupervised context [19, 20]. Other DR techniques, such as Kernel-PCA [21], scale with the number of instances. This scalability challenge makes it difficult to apply them effectively for streaming data analysis. An eventual approach for solving this issue is to control the number of instances in each chunk, but it will increase the hyperparameters of the model [22]. Self-organizing Map (SOM). Also referred to as Kohonen Neural Network is a bio-inspired method, which combines concepts from Hebbian learning, vector quantization, and competitive learning [23, 24]. Often, real-world data have important redundancies and intrinsic correlations between the correlated variables. SOMs are useful because they convert complex relationships between high-dimensional data into simple geometric relationships on a regular lattice (most often, a two-dimensional grid) [25]. SOMs methodology was introduced as a particular case of a Neural Network model. The method is a simple non-linear parametric mapping composed of a multi-layered network with Gaussian activation functions. The canonical case can be seen as a two-layered network, with the second layer generally referred to as the competition layer or feature map. Each neuron in the feature map is characterized by a reference weight vector, that has the same dimension as the input space. In spite of its simplicity, the SOM algorithm is useful for DR and visualization of high-dimensional data [26]. In addition, the method works well for unsupervised problems, and has the property of preserving the most important topological features of the reference data [24, 27, 23]. The network architecture is different from the classic feedforward case. Here, neurons are arranged on a grid. For each neuron $i$ in the feature map, there is a weight vector $\mathbf{w}_{i}^{(t)}=(w_{i1}^{(t)},\ldots,w_{id}^{(t)})$ that connects an input pattern ${\mathbf{x}}^{(t)}$ with neuron $i$. The weights are adjusted using a physiological interpretation that considers lateral neural inhibition [28, 29, 27]. For simplicity, we have here omitted the reference index for the chunk. The algorithm is iterative, composed of two main phases. The first phase is a competitive learning procedure, that has the goal of finding the neuron that better represents the current input pattern. The notion of representation is defined by the concept of closeness among multidimensional points, where all the features of the input data have the same relevance. The representative neuron is the unit in the feature map with associated weight vector closest to the current input sample. At each time $t$, in the competitive learning step, an input pattern ${\mathbf{x}}^{(t)}$ is presented to the network and a competition between neurons determines the representative neuron ${v}^{(t)}$, defined by
@@
$\Delta\mathbf{w}_{i}^{(t+1)}=\eta^{(t)}{h_{v}^{(t)}(i)}{\lVert{\mathbf{x}}^{(t)}-\mathbf{w}_{i}^{(t)}\rVert}.$ (2)
In this rule, $\eta(t)\in[0,1]$ is the learning rate [30] and the function $h_{v}^{(t)}(\cdot)$ is a smoother neighborhood kernel of the representative neuron $v$ [23]. Neighborhood functions most often are chosen from the exponential family. We include the notation for time in the neighborhood function, because a common practice is to decrease the diameter of the Gaussian with the number of iterations. Its role is to weight the region of a local neighborhood centered around the representative neuron $v$, and to control the radius of these balls. A typical choice is the Gaussian radial basis function $h_{v}(i)=\exp\{-{\lVert r_{v}-r_{i}\rVert}/2\sigma^{2}\}$, where $r_{v}$ and $r_{i}$ denote the location vectors of neurons $v$ and $i$, and $\sigma$ defines the width of the kernel. For convergence of the regression step, the limit of the kernel $h_{v}(i)$ is zero with the number of iterations [29]. Another widely selected option for a neighborhood function is the Difference of Gaussians function [31, 32].
@@
$\Delta\mathbf{w}_{i}^{(t+1)}=\eta^{(t)}h_{v}^{(t)}(i){\lVert{\mathbf{x}}^{(t)}-\mathbf{w}_{v}^{(t)}\rVert},$ (3)
where $v$ is a representative neuron in the feature map, and $h_{v}^{(t)}(\cdot)$ is the transfer function (often the same Difference of Gaussians function as in SOMs). Note that, according to Expression (3), the representative neuron has a direct effect over the other weights.
@@ 3.3 Embedding procedure for the distance matrix
To overcome these difficulties, we decided to use a less rigid projection that is coordinate-free and also that contains more information from the original space. Instead of projecting over two dimensions, we propose a reduction to a latent space in ${p}^{2}$ dimensions, where ${p}$ is the number of neurons in the feature map. Hence, our focus is on the geometric properties of the latent space. We denote by $\phi:\raisebox{2.0pt}{$\chi$}\rightarrow\mathcal{H}$ the projection function made by the self-organizing clustering method, where $\mathcal{H}$ denotes the latent space with ${p}^{2}$ dimensions. Once the model is trained, we construct a matrix $D$, where the element in position $(i,j)$ is the distance between the input data and the neuron located at $(i,j)$ position on the lattice grid. In other words, $d_{ij}=\lVert\phi({\mathbf{x}})-\mathbf{w}_{k}\rVert_{2}$, where $\mathbf{w}_{k}$ is the weight vector associated with the $(i,j)$-neuron. For computing distances among low-dimensional points in the latent space, we use the Euclidean metric. In spite of the previous reduction in dimension to value ${p}^{2}$, this is in general still large. Note that, standard empirical experiments use a squared grid of few hundred of neurons [24].
@@
With the nonlinear projection of the topographic map, the dimension is reduced from $d$ to ${p}^{2}$, which may still be too large for estimating densities. Then, we apply an additional dimensionality reduction step by moving to the first moments of the distribution represented in $D$. The first moments of a random sequence capture different aspects of the probability distribution that are useful to approximate it [38]. To evaluate the relevance of the moments in our problem, we empirically analyzed the impact of the first four moments. Let us denote $\mathbf{m}=(m_{1},m_{2},m_{3},m_{4})\in\mathbb{R}^{4}$ these moments, where $m_{1}$ is the mean, $m_{2}$ the variance, $m_{3}$ the skewness, and $m_{4}$ the kurtosis. Figure 2 presents an example of the information provided by each of the first four moments. The analysis was made offline, comparing the samples from the MNIST benchmark data versus the adversarial MNIST samples (for a description of the benchmark data see Section 4.1). Figure 2 has four plots, each having information about the mean, variance, skewness and kurtosis. Each curve represents the estimated pmf of the sequence $\mathbf{m}_{i}^{(1)},\ldots,\mathbf{m}_{i}^{(K)}$, for all chunks $S_{i}$. According to this chunk-based analysis, it is visible that the left plot in Figure 2 distinctly reveals two families of densities. One group of curves represents the pmf calculated using the original MNIST samples, while the other group corresponds to the pmf computed with the adversarial MNIST samples. This shows that the mean captures sufficient information from matrix $D$ to be able to distinguish groups of data distributions. Therefore, the proposed framework includes a final dimensionality reduction step, which involves transforming the entire matrix $D$ into a univariate sequence. Then, we represent $D$ using the first moment (mean function) of its ${p}^{2}$ elements (seen as ${p}^{2}$ independent samples of an auxiliary random variable).

### v2

@@ Abstract
Modeling non-stationary data is a challenging problem in the field of continual learning, and data distribution shifts may result in negative consequences on the performance of a machine learning model. Classic learning tools are often vulnerable to perturbations of the input covariates, and are sensitive to outliers and noise, and some tools are based on rigid algebraic assumptions. Distribution shifts are frequently occurring due to changes in raw materials for production, seasonality, a different user base, or even adversarial attacks. Therefore, there is a need for more effective distribution shift detection techniques. In this article, we introduce a self-organized continual learning framework designed for monitoring and detecting distribution changes. We explore the problem in a latent space generated by a bio-inspired self-organizing clustering and statistical aspects of the latent space. In particular, we investigate the projections made by two topology-preserving maps: the Self-Organizing Map and the Scale Invariant Map. Our method can be applied in both a supervised and an unsupervised context. We construct the assessment of changes in the data distribution as a comparison of Gaussian signals, making the proposed method fast and robust. We compare it to other unsupervised techniques, specifically Principal Component Analysis (PCA) and Kernel-PCA. Our comparison involves conducting experiments using sequences of images (based on MNIST and injected shifts with adversarial samples), chemical sensor measurements, and the environmental variable related to ozone levels. The empirical study reveals the potential of the proposed approach.
@@ 2.2 Dimensionality reduction using self-organizing clustering methods
A common approach for analyzing high-dimensional data is to apply dimensionality reduction (DR) using linear or random projections. However, both projections to the marginals and PCA present limitations for capturing the relevant structure of the original data that can cause data shift [1, 18]. On the contrary, some authors recognize the benefits of using PCA as the DR technique in a multivariate unsupervised context [19, 20]. Other DR techniques, such as Kernel-PCA [21], scale with the number of instances. This scalability challenge makes it difficult to apply them effectively for streaming data analysis. An eventual approach for solving this issue is to control the number of instances in each chunk, but it will increase the hyperparameters of the model [22]. Self-organizing Map (SOM). Also referred to as Kohonen Neural Network is a bio-inspired method, which combines concepts from Hebbian learning, vector quantization, and competitive learning [23, 24]. Often, real-world data have important redundancies and intrinsic correlations between the correlated variables. SOMs are useful because they convert complex relationships between high-dimensional data into simple geometric relationships on a regular lattice (most often, a two-dimensional grid) [25]. SOMs methodology was introduced as a particular case of a Neural Network model. The method is a simple non-linear parametric mapping composed of a multi-layered network with Gaussian activation functions. The canonical case can be seen as a two-layered network, with the second layer generally referred to as the competition layer or feature map. Each neuron in the feature map is characterized by a reference weight vector, that has the same dimension as the input space. In spite of its simplicity, the SOM algorithm is useful for DR and visualization of high-dimensional data [26]. In addition, the method works well for unsupervised problems, and has the property of preserving the most important topological features of the reference data [24, 27, 23]. The network architecture is different from the classic feedforward case. Here, neurons are arranged on a grid. Each neuron $i$ in the feature map has associated with a location in the grid that we denote to as $r_{i}$ and a weight vector $\mathbf{w}_{i}^{(t)}=(w_{i1}^{(t)},\ldots,w_{id}^{(t)})$. The weight vector has the dimension of the input space and connects an input pattern ${\mathbf{x}}^{(t)}$ with neuron $i$. The weights are adjusted using a physiological interpretation that considers lateral neural inhibition [28, 29, 27]. For simplicity, we have here omitted the reference index for the chunk. The algorithm is iterative, composed of two main phases. The first phase is a competitive learning procedure, that has the goal of finding the neuron that better represents the current input pattern. The notion of representation is defined by the concept of closeness among multidimensional points, where all the features of the input data have the same relevance. The representative neuron is the unit in the feature map with associated weight vector closest to the current input sample. At each time $t$, in the competitive learning step, an input pattern ${\mathbf{x}}^{(t)}$ is presented to the network and a competition between neurons determines the representative neuron ${v}^{(t)}$, defined by
@@
$\Delta\mathbf{w}_{i}^{(t+1)}=\eta^{(t)}h(\lVert r_{v^{(t)}}-r_{i}\rVert)\bigl{(}{\mathbf{x}}^{(t)}-\mathbf{w}_{i}^{(t)}\bigr{)}.$ (2)
In this rule, $\eta(t)\in[0,1]$ is the learning rate [30] at time $t$ and the function $h(\cdot)$ is a smoother neighborhood function. Neighborhood functions are most often chosen from the exponential family. A typical choice is $h(z)=\exp\bigl{(}-z/(2\sigma^{2})\bigr{)}$. Often, we use $\sigma=\sigma^{(t)}$ with this function decreasing when $t$ increases. Its role is to weight the region of a local neighborhood centered around the representative neuron, and to control the radius of these regions. Observe that the distance $\lVert r_{v^{(t)}}-r_{i}\rVert$ is defined in the grid (e.g. Manhattan). Another widely selected option for a neighborhood function is the Difference of Gaussians function [31, 32].
@@
$\Delta\mathbf{w}_{i}^{(t+1)}=\eta^{(t)}h(\lVert r_{v^{(t)}}-r_{i}\rVert)\bigl{(}{\mathbf{x}}^{(t)}-\mathbf{w}_{v^{(t)}}^{(t)}\bigr{)},$ (3)
where $v^{(t)}$ is a representative neuron in the feature map, and $h(\cdot)$ is the transfer function (often the same Difference of Gaussians function as in SOMs). Note that, according to Expression (3), the representative neuron has a direct effect over the other weights.
@@ 3.3 Embedding procedure for the distance matrix
To overcome these difficulties, we decided to use a less rigid projection that is coordinate-free and also that contains more information from the original space. Instead of projecting over two dimensions, we propose a reduction to a latent space in ${p}^{2}$ dimensions, where ${p}$ is the number of neurons in the feature map. Hence, our focus is on the geometric properties of the latent space. We use the way the lattice grid is defined, where each neuron in the grid is characterized by its weight vector and the location in the grid. Once the model is trained, we construct a matrix $D$, where the element in the position $(i,j)$ is the distance between the input data and the weights of the neuron located at the position $(i,j)$. In other words, $d_{i,j}=\lVert{\mathbf{x}}-\mathbf{w}_{k}\rVert_{2}$, where $\mathbf{w}_{k}$ is the weight vector associated with the $(i,j)$-neuron. We encode the information in matrix $D$ with dimensions ${p}^{2}$, that is still large. Note that, standard empirical experiments use a squared grid of few hundred of neurons [24]. With the nonlinear projection of the topographic map, the dimension is reduced from $d$ to ${p}^{2}$, which may still be too large for estimating densities. Then, we apply an additional dimensionality reduction step by moving to the first moments of the distribution represented in $D$. The first moments of a random sequence capture different aspects of the probability distribution that are useful to approximate it [38]. To evaluate the relevance of the moments in our problem, we empirically analyzed the impact of the first four moments. Let us denote $\mathbf{m}=(m_{1},m_{2},m_{3},m_{4})\in\mathbb{R}^{4}$ these moments, where $m_{1}$ is the mean, $m_{2}$ the variance, $m_{3}$ the skewness, and $m_{4}$ the kurtosis. Figure 2 presents an example of the information provided by each of the first four moments. The analysis was made offline, comparing the samples from the MNIST benchmark data versus the adversarial MNIST samples (for a description of the benchmark data see Section 4.1). Figure 2 has four plots, each having information about the mean, variance, skewness and kurtosis. Each curve represents the estimated pmf of the sequence $\mathbf{m}_{i}^{(1)},\ldots,\mathbf{m}_{i}^{(K)}$, for all chunks $S_{i}$. According to this chunk-based analysis, it is visible that the left plot in Figure 2 distinctly reveals two families of densities. One group of curves represents the pmf calculated using the original MNIST samples, while the other group corresponds to the pmf computed with the adversarial MNIST samples. This shows that the mean captures sufficient information from matrix $D$ to be able to distinguish groups of data distributions. Therefore, the proposed framework includes a final dimensionality reduction step, which involves transforming the entire matrix $D$ into a univariate sequence. Then, we represent $D$ using the first moment (mean function) of its ${p}^{2}$ elements (seen as ${p}^{2}$ independent samples of an auxiliary random variable).